OpenStudy (anonymous):

3x²+x+4=0 Help Please !

OpenStudy (zepp):

By using the quadratic formula, you will be able to find x.

OpenStudy (zepp):

\(\large \frac{-b \pm \sqrt{b^2 -4ac}}{2a}\)

OpenStudy (anonymous):

it has no roots

OpenStudy (zepp):

In this case, a = 3 b = 1 c = 4 Plug it in the formula \(\huge \frac{-(1) \pm \sqrt{(1)^2-4(3)(4)}}{2(3)}\) Then you'll find out that the discriminant is a negative, hence it has no roots; it also means that the parabola doesn't lay on the x-axis.

OpenStudy (anonymous):

you will end up with 1/12 (-i+sqrt(47)-6 i x) (i+sqrt(47)+6 i x) as your factored form

OpenStudy (anonymous):

\[1/12 (-i+\sqrt(47)-6 i x) (i+\sqrt(47)+6 i x)\]

OpenStudy (anonymous):

how did yu get that ?

OpenStudy (anonymous):

using the quadratic formula just like zepp said you get a negative under the radical

OpenStudy (anonymous):

ooh i get it ! thanks so much

OpenStudy (anonymous):

to you both !

OpenStudy (precal):

 if you look at the graph, notice that the function does not cross the x axis you will have 2 complex solutions when you use the quadratic formula, there will be a negative number under the square root sign

OpenStudy (anonymous):

\[\Delta = b^2 - 4 a c= 1^2 - 4(3)c= -47 = i^2 47\\ \sqrt \Delta = \pm i \sqrt {47}\\ x= \frac { -1 \pm i \sqrt {47}} { 6} \]